OpenStudy (anonymous):

Solve the system by substitution. -x-y-z=-8 -4x+4y+5z=7 2x+2z=4 I need step by step help please!

hero (hero):

First we should simplify each equation if possible. The first equation cn be simplified if we divide both sides by -1 The third equation can be simplified if we divide both sides by 2. The resulting system we get is: x + y + z = 8 -4x + 4y + 5 = 7 x + z = 2

hero (hero):

Next what we will do is pair the 1st and 3rd equation. Then pair the 1st and and 2nd equation to create two systems: x + y + z = 8 x + z = 2 x + y + z = 8 -4x + 4y + 5 = 7 Are you with me so far?

OpenStudy (anonymous):

Yes. So then we add them together right? @Hero

hero (hero):

Well, hang on. We may need to add or subtract depending on how things are set up. What we are looking for is to eliminate variables

hero (hero):

Now in this case, what we can do for the 1st system is subtract the 3rd equation from the first. If we do that we will get y = 6 Let me know if you understand what I did there.

OpenStudy (anonymous):

Yes I understand

hero (hero):

Okay, so we already have y = 6 Let's insert y = 6 in to the second pair. If we do that we end up with x + 6 + z = 8 -4x + 4(6) + 5z = 7

hero (hero):

Okay now after that we will end up with x + 6 + z = 8 -4x + 24 + 5z = 7

hero (hero):

And then we will subtract 6 from both sides of the top equation And subtract 24 from both sides of the bottom equation

hero (hero):

We will end up with x + z = 2 -4x + 5z = -17

hero (hero):

So now we have a system of two equations which is relatively easy to solve.

OpenStudy (anonymous):

okay

hero (hero):

From here, we can simply multiply both sides of the top equation by 4 to get 4x + 4z = 8 -4x + 5z = -17

hero (hero):

Now noticing that 4x and -4x are opposites, we can combine these equations together to eliminate x variable. When we do that we get: 9z = -9

hero (hero):

And then dividing both sides by 9 we get: z = -1

hero (hero):

Let's use the simplest possible equation to find x.

hero (hero):

x + z = 2

hero (hero):

Since z = -1 we have x - 1 = 2

hero (hero):

So what does x equal?

OpenStudy (anonymous):

3?

hero (hero):

Correct. So what are the cooridinates of the solution point (x, y, z)?

OpenStudy (anonymous):

3,6,-1

hero (hero):

Yes correct.

hero (hero):

The system is correct, but I made some errors above. There are places where I wrote 5 instead of 5z. I hope you recognize it.

hero (hero):

Anyway, I will proceed to the other question.

OpenStudy (anonymous):

yeah I saw it that you!